# Unified description of sum rules and duality between CP phases and unitarity triangles through third-order rephasing invariants

### v1

@@ III Third-order rephasing invariants with determinant
$\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.16^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}-\begin{pmatrix}1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ \end{pmatrix}+2\begin{pmatrix}1.052^{\circ}&0.035^{\circ}&92.40^{\circ}\\ 0.035^{\circ}&92.40^{\circ}&1.052^{\circ}\\ 92.40^{\circ}&1.052^{\circ}&0.035^{\circ}\\ \end{pmatrix}$ (48)
$=\begin{pmatrix}93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ 93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ 93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ \end{pmatrix}=\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ \end{pmatrix}\Delta-\Pi.$ (49)
This equation precisely reveals the underlying order between scattered and chaotic values of the CP phases, angles, and third-order invariants.
@@
$\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.16^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}+\begin{pmatrix}1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ \end{pmatrix}+2\begin{pmatrix}-0.0019^{\circ}&-22.45^{\circ}&-64.06^{\circ}\\ -64.06^{\circ}&-0.0019^{\circ}&-22.45^{\circ}\\ -22.45^{\circ}&-64.06^{\circ}&-0.0019^{\circ}\\ \end{pmatrix}$ (52)
@@
The reason why the expression is written in terms of simple sums and differences is that the angle matrix $\Phi$ and the phase matrix $\Delta$ are defined so as not to share any common matrix elements of $V$.
@@ IV Summary
In this letter, we demonstrate that all nine angles of the unitarity triangles and CP phases in the nine different Euler-angle-like representations are expressed by products of third-order rephasing invariants $V_{\alpha i}V_{\beta j}V_{\gamma k}/\det V$. The fourth-order angles and the fifth-order phases are composed of two and three invariants. The sum rules relating these nine angles and phases are also decomposed into terms of the third-order invariants.
@@
These results indicate that the third-order invariants play an essential role in understanding of flavored CP violation. Similarly, sixth-, seventh-, and higher-order invariants can also be constructed, and simple sum rules and dualities are expected among their arguments and lower-order invariants. The framework developed in this study complements the traditional Jarlskog invariants, offering a deeper structural understanding of CP violation. Moreover, this approach is not limited to the CKM matrix but can also be applied to the MNS matrix and extended flavor structures, providing an effective analytical tool for precision measurements and new physics searches.

### v2

@@ III Third-order rephasing invariants with determinant
$\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.15^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}-\begin{pmatrix}1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ \end{pmatrix}+2\begin{pmatrix}1.052^{\circ}&0.035^{\circ}&92.40^{\circ}\\ 0.035^{\circ}&92.40^{\circ}&1.052^{\circ}\\ 92.40^{\circ}&1.052^{\circ}&0.035^{\circ}\\ \end{pmatrix}$
$=\begin{pmatrix}93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ 93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ 93.49^{\circ}&93.49^{\circ}&93.49^{\circ}\\ \end{pmatrix}=\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ \end{pmatrix}\Delta-\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\\ \end{pmatrix}\pi.$ (46)
The reason why the expression is written by the simple difference is that the angle matrix $\Phi$ and the phase matrix $\Delta$ are defined so as not to share any common matrix elements of $V$.
@@
$\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.15^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}+\begin{pmatrix}1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ \end{pmatrix}+2\begin{pmatrix}-0.0019^{\circ}&-22.45^{\circ}&-64.06^{\circ}\\ -64.06^{\circ}&-0.0019^{\circ}&-22.45^{\circ}\\ -22.45^{\circ}&-64.06^{\circ}&-0.0019^{\circ}\\ \end{pmatrix}$
@@
These equations precisely reveal the underlying order between scattered and chaotic values of the CP phases, angles, and third-order invariants.
@@ IV Summary
In this paper, we demonstrate that all nine angles of the unitarity triangles and CP phases in the nine different Euler-angle-like representations are expressed by products of third-order rephasing invariants $V_{\alpha i}V_{\beta j}V_{\gamma k}/\det V$. The fourth-order angles and the fifth-order phases are composed of two and three invariants. This fact shows that the fifth-order invariants are not genuinely irreducible. The sum rules relating these nine angles and phases are also decomposed into terms of the third-order invariants.
@@
These results indicate that the third-order invariants play an essential role in the understanding of flavored CP violation. Similarly, sixth-, seventh-, and higher-order invariants can also be constructed, and simple sum rules and dualities are expected among their arguments and lower-order invariants. The framework developed in this study complements the traditional Jarlskog invariants, offering a deeper structural understanding of CP violation. Moreover, this approach is not limited to the CKM matrix but can also be applied to the MNS matrix, because varying the values of the Majorana phases corresponds to rephasing transformations of the neutrinos and the value of $\delta$ remains unaffected. Therefore, this methodology provides an effective analytical tool for precision measurements and new physics searches.